Suppose you want to have $500,000 by the end of 7 years. You decide to deposit an equal amount of money each month (starting next month) into a bank account that advertises a nominal rate of 4% compounded monthly. How much should you deposit each month?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A B. FV I 2 3 Amount required at end 7 years APR Monthly interest rate No of monthly deposit in 7 years $500,000.00 4% 0.3333

Cell reference -

A В. FV 3 Amount required at end 7 years APR Monthly interest rate No of monthly deposit in 7 years 500000 0.04 =C3/12 =7*12

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.